# Black Hole Entropy, Quantum Corrections and EFT Transitions

### v1

@@ Abstract
We revisit and study quantum corrections to the supersymmetric entropy of BPS black holes in 4d $\mathcal{N}=2$ effective field theories (EFTs), which can be obtained from Type IIA string theory compactified on a Calabi–Yau threefold. Macroscopically, these corrections arise from an infinite series of higher-derivative F-terms that encode certain modifications to the two-derivative supergravity effective action. Within the large volume regime, we analyze in detail the moduli dependence of these semi-classical contributions and explore their implications for the black hole entropy. As a byproduct, we show that the entropy captures, in a rather intricate way, the transition between four- and five-dimensional dual EFT descriptions. In fact, the expansion parameter $\alpha$ controlling the relevant asymptotic series can be related to the ratio of the black hole horizon and the Kaluza-Klein scale, given here by the inverse D0-brane mass. Furthermore, we are able to resum the series into a well-behaved convergent expression for all values of $\alpha$. This demonstrates, in turn, that (stable) black holes can, indeed, probe scales besides the quantum gravity cutoff. More precisely, by examining two representative BPS systems —the D0-D2-D4 and D2-D6 black hole solutions— we explicitly illustrate how highly non-local (perturbative) quantum effects resolve the divergences, ultimately leading to a well-defined entropy function. Additionally, in certain cases, we show that one can take a suitable decompactification limit to 5d and verify that the corrected entropy function reproduces the exact microstate counting of the underlying five-dimensional black string. Our results also clarify the role of non-perturbative quantum corrections, which, remarkably, do not modify any of our prior conclusions.
@@ 1 Introduction and Summary
As is widely expected —given the non-renormalizability of Einstein’s gravity at the quantum level, the structure of any gravitational Effective Field Theory (EFT) should be such that the suppression of generic higher-derivative and higher-curvature operators relative to the Einstein-Hilbert term is dictated by some specific energy scale vandeHeisteeg:2022btw ; Cribiori:2022nke ; vandeHeisteeg:2023ubh ; vandeHeisteeg:2023dlw ; Castellano:2023aum , namely the quantum gravity or species cut-off Dvali:2007hz ; Dvali:2009ks ; Dvali:2010vm ; Dvali:2012uq (see also Castellano:2024bna for an comprehensive treatment of this subject). However, it is also known that the same kind of higher-dimensional operators in EFTs coupled to gravity —e.g., those arising from string or M-theory— oftentimes exhibit an explicit suppression by (even parametrically) lower scales, such as the Kaluza-Klein (KK) mass Castellano:2023aum ; Aoufia:2024awo ; Calderon-Infante:2025ldq . In fact, a simple realization of this scenario is given by four-dimensional $\mathcal{N}=2$ supergravities obtained from compactifying Type IIA string theory on a Calabi–Yau threefold. There, the lightest KK scale can sometimes correspond to the D0-brane mass, whose associated species cutoff is given by certain 5d Planck scale Castellano:2022bvr . What happens then, when we reach the KK scale, is that the 4d EFT description breaks down, signaling that we should switch to the dual five-dimensional EFT arising from considering M-theory on the same Calabi–Yau space. An interesting question that one can ask, given this state of affairs, is whether and how such EFT transitions could be characterized using the thermodynamics of existing black hole solutions in the theory. Indeed, one may expect that, in general, the solutions themselves must be subject to some kind of phase transition. For instance, these could correspond to transitions of the Gregory-Laflamme Gregory:1993vy ; Gregory:1994bj or Horowitz-Polchinski Horowitz:1996nw ; Horowitz:1997jc type. [^1] See Brustein:2021ifl ; Chen:2021emg ; Chen:2021dsw ; Urbach:2022xzw ; Balthazar:2022szl ; Balthazar:2022hno ; Ceplak:2023afb ; Herraez:2024kux ; Albertini:2024hwi ; Chu:2024ggi ; Emparan:2024mbp ; Ceplak:2024dxm ; Bedroya:2024igb ; Chu:2025fko for recent developments regarding this kind of transitions in quantum gravity and string theory. Nevertheless, by sticking to stable BPS black holes it might be possible that some of these solutions are actually not subject to any such instability, and that the associated thermodynamic quantities smoothly interpolate between possibly very different EFT regimes. This last possibility is particularly compelling because it would mean that for certain non-perturbative gravitational objects in the theory, we would be able to describe in detail their behavior within the transition regime. Also, this means that one could define a family of solutions which explicitly interpolates and glues between two complementary EFT descriptions living in different number of spacetime dimensions. The goal of the present work is to explore this possibility.
@@
For this purpose, we investigate the behavior of a restricted set of quantum corrections to the black hole entropy in 4d $\mathcal{N}=2$ supersymmetric effective field theories, focusing on the convergence properties of the associated perturbative expansions. More concretely, the black hole solutions we consider are BPS configurations Bogomolny:1975de ; Prasad:1975kr , and to them we can associate some entropy, $\mathcal{S}_{\rm BH}$, that can be determined solely as a function of its gauge charges LopesCardoso:1998tkj ; LopesCardoso:1999cv ; LopesCardoso:1999fsj ; LopesCardoso:1999xn , and which is moreover protected by supersymmetry. The main reason for choosing this particular set-up is that, according to the literature (see, e.g., Mohaupt:2000mj ; Ooguri:2004zv and references therein), it is strongly believed that all relevant corrections to the aforementioned indexed quantity are already captured by the low energy (supergravity) EFT in the form of an infinite number of higher-dimensional and higher-curvature local BPS operators involving the (anti-self-dual parts of the) graviton and graviphoton field strengths. Such operators contribute non-trivially to the entropy of supersymmetric black holes Wald:1993nt ; Iyer:1994ys , and in fact can be seen to exhibit interesting asymptotic behaviors for certain values of the black hole charges Cribiori:2022nke ; Cribiori:2023ffn ; Calderon-Infante:2023uhz ; Basile:2023blg ; Basile:2024dqq ; Bedroya:2024ubj ; Herraez:2024kux ; smallBHsMunich . In this work, we show that the quantum corrections to the entropy organize themselves into an asymptotic series whose expansion parameter $\alpha$ is related to the ratio between the M-theory circle radius —computed as the inverse D0-brane mass, and the size of the black hole horizon. We therefore identify a transition regime corresponding to $\alpha\sim\mathcal{O}(1)$, which is equivalent to consider black hole solutions whose radius becomes comparable to that of the M-theory circle. Moreover, for all values of $\alpha$, we are able to perform a resummation of the underlying asymptotic series, ultimately providing explicit and convergent expressions.
@@
To illustrate this point, we analyze in detail two sub-classes of BPS solutions of the attractor mechanism close to the large radius point, namely D0-D2-D4 brane configurations, and systems exhibiting only D2- and D6-brane charges. In both cases, the highly non-local perturbative effects induced by the infinite tower of light Kaluza-Klein states are crucial to cancel the UV divergences exhibited by the four-dimensional EFT, and they in turn allow us to resum the asymptotic series in an exact manner —following the same strategy as in Gopakumar:1998ii ; Gopakumar:1998jq . Interestingly, we can also evaluate the most dominant non-perturbative effects, but they do not seem to play a major role neither in the construction of the solutions solving the attractor equations, nor in the cancellation of the divergences of the entropy. In fact, the two examples analyzed in this work turn out being somewhat complementary. Indeed, from a computational perspective, the resummation procedure that we use for the D0-D2-D4 works as long as the black hole carries D0-brane charge. The study of the D2-D6 configuration is therefore crucial to explain how to extend our prescription to a more general choice of charges.
@@
In the case of the D0-D2-D4 configuration we end up with a quantum-corrected entropy that is well-defined for all values of the parameter $\alpha$. In particular, we show that for $\alpha\sim\mathcal{O}(1)$ the asymptotic series stops being valid (for any of its finite-order truncations), thereby reflecting that the four-dimensional EFT breaks down. Therefore, upon crossing this transition regime, the solution itself should be most naturally regarded as a five-dimensional black string wrapped on the extra circle. Moreover, in the limit of large $\alpha$, we recover an infinitely extended black string living in five non-compact dimensions. Crucially, we show how the resummed version of the higher-derivative corrections to the entropy get diluted —except for one particular term— so as to precisely reproduce the exact microstate counting of the five-dimensional black string, which is also in agreement with other macroscopic computations in 5d $\mathcal{N}=1$ supergravity. This provides, in turn, a highly non-trivial check of our result. On the other hand, for the D2-D6 configuration, even if we are able to describe analytically the transition regime, the five-dimensional uplift of the 4d black hole crucially carries Kaluza–Klein monopole charge, and it exists only as long as there is a $\mathbf{S}^{1}$ direction in the theory. In this second case, we are still capable to glue two different EFT descriptions but we cannot explore the purely five-dimensional regime (i.e., the large $|\alpha|$ limit).
@@ 2 Review: BPS Black Holes in Four Dimensions
with $A,B=0,1,\ldots,h^{1,1}(X_{3})$. We denote by $z^{a}=b^{a}+it^{a}$, $a=1,\ldots,h^{1,1}(X_{3})$, the scalar fields describing the complexified Kähler (or vector multiplet) moduli space of the theory, whereas $\xi^{p}$, $p=1,\ldots,h^{2,1}(X_{3})+1$, belong to the hypermultiplets instead. The field strengths $F^{B}=dA^{B}$ correspond to $U(1)$ gauge bosons normalized so that they have integrally-quantized charges. In this work we will restrict ourselves to the vector multiplet sector, since the black hole solutions we are most interested in here only depend on the latter.
@@
given a local patch where $X^{0}$ does not vanish anywhere. This also implies that the entire geometry of the vector multiplet moduli space can be encoded into a holomorphic function $\mathcal{F}(X^{A})$, usually referred to as the prepotential Craps:1997gp ; Craps:1997nv . This function is moreover homogeneous of degree two, meaning that it satisfies $\mathcal{F}=\frac{1}{2}X^{A}\mathcal{F}_{A}$, where $\mathcal{F}_{A}=\partial_{X^{A}}\mathcal{F}$.
@@
On the other hand, these theories are known to present —beyond the two-derivative lagrangian (2.1)— interesting higher-dimensional and higher-curvature corrections. Some of these terms are furthermore $\frac{1}{2}$-BPS which means, in practice, that they are protected by supersymmetry from receiving certain quantum corrections. This implies, in turn, that their dependence with respect to the moduli fields $z^{a}$ can be sometimes determined exactly. Using standard $\mathcal{N}=2$ superspace notation, they can be written as follows Bershadsky:1993ta ; Bershadsky:1993cx ; Antoniadis:1993ze ; Antoniadis:1995zn  [^2] Note that the $g=0$ contribution gives precisely the prepotential term in $\mathcal{N}=2$ supergravity, upon identifying $\mathcal{F}_{0}(\mathcal{X}^{A})\equiv\mathcal{F}(\mathcal{X}^{A})$ as functions of the chiral superfields (2.9).
[…]
@@
[…]
with $X^{A}$ denoting the bottom (i.e., scalar) components of the reduced chiral superfields deRoo:1980mm
$\mathcal{X}^{A}=X^{A}+\frac{1}{2}\epsilon_{ij}\theta^{i}\sigma^{\mu\nu}\theta^{j}\left(F^{A,-}_{\mu\nu}-\frac{i}{2}e^{K/2}\bar{X}^{A}W_{\mu\nu}^{-}\right)+\ldots\,,$ (2.9)
and where the precise index contractions appearing in (2.8) can be deduced from eqs. (2.5) and (2.6). Let us remark that not all the purely bosonic terms that can be extracted from the superspace lagrangian (2.5) are quadratic in the Riemann tensor. In fact, if instead of using the $\theta^{2}$-component of $\mathcal{W}^{ij}_{\mu\nu}$ we rather insert the maximal $\theta^{4}$-term —which contains a piece proportional to the anti-self dual component of the antisymmetric tensor $\nabla_{[\mu}\nabla^{\sigma}W^{kl,-}_{\sigma|\nu]}$, we obtain a local operator in the action $W^{ij,-}_{\mu\rho}\nabla^{\rho}\nabla^{\sigma}W^{kl,-}_{\sigma\nu}\epsilon_{ik}\epsilon_{jl}$ that is quadratic in the graviphoton field strength and moreover contains two covariant derivatives Bergshoeff:1980is ; LopesCardoso:1998tkj . Such a term would then be linear in the Riemann tensor, and in fact turns out being the only one contributing to the entropy at the four-derivative level LopesCardoso:1998tkj (see discussion in Section 2.2 below).
Interestingly, as originally noticed in Gopakumar:1998ii ; Gopakumar:1998jq , one can compute all perturbative and non-perturbative stringy $\alpha^{\prime}$-corrections in $\mathcal{F}_{g}(X^{A})$ for $g\geq 0$ using the duality between Type IIA string theory on $X_{3}$ and M-theory compactified on $X_{3}\times\mathbf{S}^{1}$. This exploits the fact that the string coupling belongs to a hypermultiplet, which is decoupled from the vector multiplets at the two-derivative level Candelas:1990pi , such that it can be freely tuned at will. Hence, for a single hypermultiplet of mass $m=|Z|$ in 4d Planck units, with $Z=e^{K/2}\left(q_{A}X^{A}-p^{A}\mathcal{F}_{A}\right)$ being its central charge, one indeed obtains a generating function via a Schwinger-like one-loop computation as follows (see, e.g., Bastianelli:2008cu ; Dunne:2004nc and references therein)

### v2

@@ Abstract
We revisit and study quantum corrections to the supersymmetric entropy of BPS black holes in 4d $\mathcal{N}=2$ effective field theories (EFTs), which can be obtained from Type IIA string theory compactified on a Calabi–Yau threefold. Macroscopically, these corrections arise from an infinite series of higher-derivative F-terms that encode certain modifications to the two-derivative supergravity effective action. Within the large volume regime, we analyze in detail the moduli dependence of these semi-classical contributions and explore their implications for the black hole entropy. As a byproduct, we show that the latter captures, in a rather intricate way, the transition between four- and five-dimensional dual EFT descriptions. In fact, the expansion parameter $\alpha$ controlling the relevant asymptotic series can be related to the ratio of the black hole horizon and the Kaluza-Klein length-scale, given here by the inverse D0-brane mass. Furthermore, we are able to resum the series into a well-behaved convergent expression for all values of $\alpha$. This demonstrates, in turn, that (stable) black holes can, indeed, probe scales besides the quantum gravity cutoff. More precisely, by examining two representative BPS systems —the D0-D2-D4 and D2-D6 black hole solutions— we explicitly illustrate how highly non-local yet perturbative quantum effects resolve the divergences, ultimately leading to a well-defined entropy function. Additionally, in special cases, we show that one can take a suitable decompactification limit to 5d and verify that the corrected entropy function reproduces the exact microstate counting of the underlying five-dimensional black string. Our results also clarify the role of certain non-perturbative quantum corrections, which, remarkably, do not modify any of our prior conclusions.
@@ 1 Introduction and Summary
As is widely expected —given the non-renormalizability of Einstein’s gravity at the quantum level, the structure of any gravitational Effective Field Theory (EFT) should be such that the suppression of generic higher-derivative and higher-curvature operators relative to the Einstein-Hilbert term is dictated by some specific energy scale vandeHeisteeg:2022btw ; Cribiori:2022nke ; vandeHeisteeg:2023ubh ; vandeHeisteeg:2023dlw ; Castellano:2023aum , namely the quantum gravity or species cut-off Dvali:2007hz ; Dvali:2009ks ; Dvali:2010vm ; Dvali:2012uq (see also Castellano:2024bna for a comprehensive treatment of this subject). However, it is also known that the same kind of higher-dimensional operators in EFTs coupled to gravity —e.g., those arising from string or M-theory— quite often exhibit an explicit suppression by (even parametrically) lower scales, such as the Kaluza-Klein (KK) mass Castellano:2023aum ; Aoufia:2024awo ; Calderon-Infante:2025ldq . In fact, a simple realization of this scenario is given by four-dimensional $\mathcal{N}=2$ supergravities obtained from compactifying Type IIA string theory on a Calabi–Yau threefold. There, the lightest KK scale can sometimes correspond to the D0-brane mass, whose associated species cutoff is given by certain 5d Planck scale Castellano:2022bvr . What happens then, when we reach the KK scale, is that the 4d description breaks down, signaling that we should switch to the dual five-dimensional EFT arising from considering M-theory on the same Calabi–Yau space. An interesting question that one can ask, given this state of affairs, is whether and how such EFT transitions could be characterized using the thermodynamics of existing black hole solutions in the theory. Indeed, one may expect that, in general, the solutions themselves must be subject to some kind of phase transition. For instance, these could correspond to transitions of the Gregory-Laflamme Gregory:1993vy ; Gregory:1994bj or Horowitz-Polchinski Horowitz:1996nw ; Horowitz:1997jc type. [^1] See Brustein:2021ifl ; Chen:2021emg ; Chen:2021dsw ; Urbach:2022xzw ; Balthazar:2022szl ; Balthazar:2022hno ; Ceplak:2023afb ; Herraez:2024kux ; Albertini:2024hwi ; Chu:2024ggi ; Emparan:2024mbp ; Ceplak:2024dxm ; Bedroya:2024igb ; Chu:2025fko for recent developments regarding this kind of transitions in quantum gravity and string theory. Nevertheless, by sticking to stable BPS black holes it might be possible that some of these solutions do not actually suffer from any such instability, and that the associated thermodynamic quantities smoothly interpolate between possibly very different EFT regimes. This last possibility is particularly compelling because it would mean that for certain non-perturbative gravitational objects in the theory, we would be able to describe in detail their behavior within the transition regime. Also, this means that one could define a family of solutions which explicitly interpolates and glues between two complementary EFT descriptions living in different number of spacetime dimensions. The goal of the present work is to explore this latter scenario.
@@
For this purpose, we investigate the behavior of a restricted set of quantum corrections to the black hole entropy in 4d $\mathcal{N}=2$ supersymmetric effective field theories, focusing on the convergence properties of their associated perturbative expansions. More concretely, the black hole solutions we consider are BPS configurations Bogomolny:1975de ; Prasad:1975kr , and to them we can associate some indexed entropy, $\mathcal{S}_{\rm BH}$, that can be determined solely as a function of its gauge charges LopesCardoso:1998tkj ; LopesCardoso:1999cv ; LopesCardoso:1999fsj ; LopesCardoso:1999xn , and which is moreover protected by supersymmetry. The main reason for choosing this particular set-up is that, according to the literature (see, e.g., Mohaupt:2000mj ; Ooguri:2004zv and references therein), it is strongly believed that all relevant corrections to the aforementioned quantity are already captured by the low energy (supergravity) EFT in the form of an infinite number of higher-dimensional and higher-curvature local BPS operators involving the (anti-self-dual parts of the) graviton and graviphoton field strengths. Such operators contribute non-trivially to the entropy of supersymmetric black holes Wald:1993nt ; Iyer:1994ys , and in fact can be seen to exhibit an interesting behavior for certain values of the black hole charges Cribiori:2022nke ; Cribiori:2023ffn ; Calderon-Infante:2023uhz ; Basile:2023blg ; Basile:2024dqq ; Bedroya:2024ubj ; Herraez:2024kux ; Calderon-Infante:2025pls . In this work, we show that the quantum corrections to the entropy organize themselves into an asymptotic series whose complex expansion parameter $\alpha$ is related to the ratio between the M-theory circle radius —computed as the inverse D0-brane mass, and the size of the black hole horizon. We therefore identify a transition regime corresponding to $|\alpha|=\mathcal{O}(1)$, which is equivalent to considering black hole solutions whose radius becomes comparable to that of the M-theory circle. Moreover, for all values of $\alpha$, we are able to perform a resummation of the underlying asymptotic series, ultimately providing explicit and convergent expressions.
@@
To illustrate this point, we analyze in detail two sub-classes of BPS solutions of the attractor mechanism close to the large radius point, namely D0-D2-D4 brane configurations, and systems exhibiting only D2- and D6-brane charges. In both cases, the highly non-local perturbative effects induced by the infinite tower of light Kaluza-Klein states are crucial to cancel the UV divergences exhibited by the four-dimensional EFT, and they in turn allow us to resum the asymptotic series in an exact manner —following the same strategy as in Gopakumar:1998ii ; Gopakumar:1998jq . Interestingly, we can also evaluate the most dominant non-perturbative effects, but they do not seem to play a major role neither in the construction of the solutions solving the attractor equations, nor in the cancellation of the divergences of the entropy. In fact, the two examples analyzed in this work turn out being somewhat complementary. Indeed, from a computational perspective, the resummation procedure that we use for the D0-D2-D4 works as long as $\alpha$ is not purely imaginary. The study of the D2-D6 configuration is therefore crucial to explain how to extend our contour prescription to a more general choice of charges.
@@
In the case of the D0-D2-D4 configuration we end up with a quantum-corrected entropy that is well-defined for all values of the now real parameter $\alpha$. In particular, we show that for $\alpha=\mathcal{O}(1)$ the asymptotic series stops being valid (for any of its finite-order truncations), thereby reflecting that the four-dimensional EFT breaks down. Therefore, upon crossing this transition regime, the solution itself should be most naturally regarded as a five-dimensional black string wrapped on the extra circle. Moreover, in the limit of large $\alpha$, we recover an infinitely extended black string living in five non-compact dimensions. Crucially, we show how the resummed version of the higher-derivative corrections to the entropy get diluted —except for one particular term— so as to precisely reproduce the exact microstate counting of the five-dimensional black string, which is also in agreement with other macroscopic computations in 5d $\mathcal{N}=1$ supergravity. This provides, in turn, a highly non-trivial check of our result. On the other hand, for the D2-D6 configuration, even if we are able to describe analytically the transition regime, the five-dimensional uplift of the 4d black hole crucially carries Kaluza–Klein monopole charge, and it exists only as long as there is a compact $\mathbf{S}^{1}$ direction in the theory. In this second case, we are still capable to glue two different EFT descriptions but we cannot explore the purely five-dimensional regime (i.e., the large $|\alpha|$ limit).
@@ 2 Review: BPS Black Holes in Four Dimensions
with $A,B=0,1,\ldots,h^{1,1}(X_{3})$. We denote by $z^{a}=b^{a}+it^{a}$, $a=1,\ldots,h^{1,1}(X_{3})$, the scalar fields describing the complexified Kähler (or vector multiplet) moduli space of the theory, whereas $\xi^{p}$, $p=1,\ldots,h^{2,1}(X_{3})+1$, belong to the hypermultiplets instead. The field strengths $F^{B}=dA^{B}$ correspond to $U(1)$ gauge bosons normalized so that they have integrally-quantized charges. In this work, we will restrict ourselves to the vector multiplet sector, since the black hole solutions we are most interested in here only depend on the latter.
@@
given a local patch where $X^{0}$ is nowhere vanishing. This also implies that the entire geometry of the vector multiplet moduli space can be encoded into a holomorphic function $\mathcal{F}(X^{A})$, usually referred to as the prepotential Craps:1997gp ; Craps:1997nv . This function is moreover homogeneous of degree two, meaning that it satisfies $\mathcal{F}=\frac{1}{2}X^{A}\mathcal{F}_{A}$, where $\mathcal{F}_{A}=\partial_{X^{A}}\mathcal{F}$.
@@
On the other hand, these theories are known to present —beyond the two-derivative Lagrangian (2.1)— interesting higher-dimensional and higher-curvature corrections. Some of these terms are furthermore $\frac{1}{2}$-BPS which means, in practice, that they are protected by supersymmetry from receiving certain quantum corrections. This implies, in turn, that their dependence with respect to the moduli fields $z^{a}$ can sometimes be determined exactly. Using standard $\mathcal{N}=2$ superspace notation, they can be written as follows Bershadsky:1993ta ; Bershadsky:1993cx ; Antoniadis:1993ze ; Antoniadis:1995zn  [^2] Note that the $g=0$ contribution gives precisely the prepotential term in $\mathcal{N}=2$ supergravity, upon identifying $\mathcal{F}_{0}(\mathcal{X}^{A})\equiv\mathcal{F}(\mathcal{X}^{A})$ as functions of the chiral superfields (2.9).
$\mathcal{L}_{\rm h.d.}\,\supset\,-\frac{i}{2}\sum_{g\geq 1}\int\text{d}^{4}\theta\,\mathcal{F}_{g}(\mathcal{X}^{A})\,\left(\mathcal{W}^{ij}\mathcal{W}_{ij}\right)^{g}\ +\ \text{h.c.}\,$, (2.5)
@@
$\mathcal{L}_{\rm h.d.}\,\supset\,-\frac{i}{2}\sum_{g\geq 1}\mathcal{F}_{g}(X^{A})\,\mathcal{R}_{-}^{2}\,W_{-}^{2g-2}\ +\ \text{h.c.}\,$, (2.8)
with $X^{A}$ denoting the bottom (i.e., scalar) constituents of the reduced chiral superfields deRoo:1980mm
$\mathcal{X}^{A}=X^{A}+\frac{1}{2}\epsilon_{ij}\theta^{i}\sigma^{\mu\nu}\theta^{j}\left(F^{A,-}_{\mu\nu}-ie^{K/2}\bar{X}^{A}W_{\mu\nu}^{-}\right)+\ldots\,$, (2.9)
and where the precise index contractions appearing in (2.8) can be deduced from eqs. (2.5) and (2.6). Let us remark that not all the purely bosonic terms that can be extracted from the superspace Lagrangian (2.5) are quadratic in the Riemann tensor. In fact, if instead of using the $\theta^{2}$-component of $\mathcal{W}^{ij}_{\mu\nu}$ we rather insert the maximal $\theta^{4}$-term —which contains a piece proportional to the anti-self dual combination of the antisymmetric tensor $\nabla_{[\mu}\nabla^{\sigma}W^{kl,-}_{\sigma|\nu]},$ we obtain a local operator in the action $W^{ij,-}_{\mu\rho}\nabla^{\rho}\nabla^{\sigma}W^{kl,-}_{\sigma\nu}\epsilon_{ik}\epsilon_{jl}$ that is quadratic in the graviphoton field strength and moreover contains two covariant derivatives Bergshoeff:1980is ; LopesCardoso:1998tkj . Such a term would then be linear in the Riemann tensor, and in fact turns out being the only one contributing to the entropy at the four-derivative level LopesCardoso:1998tkj (see discussion in Section 2.2 below).
Interestingly, as originally noticed in Gopakumar:1998ii ; Gopakumar:1998jq , one can compute all perturbative and non-perturbative stringy $\alpha^{\prime}$-corrections in $\mathcal{F}_{g}(X^{A})$ for $g\geq 0$ using the duality between Type IIA string theory on $X_{3}$ and M-theory compactified on $X_{3}\times\mathbf{S}^{1}$. This exploits the fact that the string coupling belongs to a hypermultiplet, which is decoupled from the vector multiplets at the two-derivative level Candelas:1990pi , such that it can be freely tuned at will. Hence, for a single hypermultiplet of mass $m=|Z|$ in 4d Planck units, with $Z=e^{K/2}\left(p^{A}\mathcal{F}_{A}-q_{A}X^{A}\right)$ being its central charge, one indeed obtains a generating function via a Schwinger-like one-loop computation as follows (see, e.g., Bastianelli:2008cu ; Dunne:2004nc and references therein)